OpenStudy (anonymous):

Find the system of equations that represents the following matrix equation. [3 2 ] [x ] = [ 9 ] 1 -2 y 11

OpenStudy (anonymous):

3x+2y=9 x-2y=11

OpenStudy (anonymous):

My choices are a. 3y + 2x = 9 1y – 2x = 11   b. 3y + 2x = 9 1x – 2y = 11   c. 3x + 2y = 9 1y – 2x =11   d. 3x + 2y = 9 1x – 2y = 11